# A Symmetric Regressor for MRI-Based Assessment of Striatal Dopamine Transporter Uptake in Parkinson’s Disease

### v1

@@ Abstract
Dopamine transporter (DAT) imaging is commonly used for monitoring Parkinson’s disease (PD), where striatal DAT uptake amount is computed to assess PD severity. However, DAT imaging has a high cost and the risk of radiance exposure and is not available in general clinics. Recently, MRI patch of the nigral region has been proposed as a safer and easier alternative. This paper proposes a symmetric regressor for predicting the DAT uptake amount from the nigral MRI patch. Acknowledging the symmetry between the right and left nigrae, the proposed regressor incorporates a paired input-output model that simultaneously predicts the DAT uptake amounts for both the right and left striata. Moreover, it employs a symmetric loss that imposes a constraint on the difference between right-to-left predictions, resembling the high correlation in DAT uptake amounts in the two lateral sides. Additionally, we propose a symmetric Monte-Carlo (MC) dropout method for providing a fruitful uncertainty estimate of the DAT uptake prediction, which utilizes the above symmetry. We evaluated the proposed approach on $734$ nigral patches, which demonstrated significantly improved performance of the symmetric regressor compared with the standard regressors while giving better explainability and feature representation. The symmetric MC dropout also gave precise uncertainty ranges with a high probability of including the true DAT uptake amounts within the range.
{IEEEkeywords}
Brain MRI, DAT imaging, DAT uptake assessment, deep learning, dopamine transporter, Monte-Carlo dropout, Parkinson’s disease, prediction interval, severity assessment, SPECT, SMWI, substantia nigra, symmetric loss, symmetric regressor, uncertainty
@@ 1 Introduction
Recent studies suggested using MRI as a less invasive and readily available alternative imaging method for monitoring PD [6, 7]. In MRI-based assessment, the loss of nigral hyperintensity is generally used for PD diagnosis as it corresponds well with the low DAT uptake. Nonetheless, a quantified severity assessment is not achievable by only manual observation of such intensity. Therefore, a computerized method for predicting the DAT uptake amount (as measured in SPECT) using the nigral MRI can be substantially useful.
In this paper, we propose a symmetric deep regressor for predicting the DAT uptake amount from the MRI patch covering the nigral region. While the typical regressor follows a one-input-one-output structure with a loss function based on the mean squared error (MSE) w.r.t. the known true output, our symmetric model adopts a paired input-output structure. Our regressor takes both the right and left nigral patches, simultaneously generating DAT uptake predictions for both the right and left striata. Additionally, we employ a symmetric loss that penalizes high differences between right and left predictions, addressing the symmetry between the two nigral regions and the strong correlation between DAT uptakes in the right and left striata. The proposed formulation facilitates explicit knowledge sharing between the two prediction tasks, thereby, allowing for enhanced feature learning. Fig. 1 illustrates the proposed symmetric regressor model.
@@
  • We propose a symmetric regressor for the striatal DAT uptake prediction, which utilizes the lateral symmetry and correlation of the right and left nigral patch inputs and the corresponding DAT uptake outputs.
@@ 2.2 DAT Uptake Prediction using MRI
The use case of MRI is not limited to only PD diagnosis. Researchers have also used the nigral status in MRI to correlate with the SBR in SPECT. Bae et al. [18] compared the nigral hyperintensity observed in 3T SWI with the SPECT SBR in Parkinsonism. In their later work [19], they used a similar SWI-to-SBR comparison approach for idiopathic rapid eye movement disorder. Both of these works revealed good agreement between the SWI observation and the SBR score. Recently, susceptibility map-weighted imaging (SMWI) [20] has been proposed for improving the image quality by providing with better contrast in neuroimaging. Nam et al. [21] utilized SMWI to enhance the visibility of nigrosome for PD diagnosis, which allowed for better identification of the hyperintensity. Bae et al. [22] also compared SWI and SMWI images of the substantia nigra in Parkinsonism, and concluded that SMWI provides improved quality.
@@ 2.3 Deep Learning Based Prediction
In this work, we utilize both the right and left nigral patches to predict SBRs of both the right and left striata. Leveraging the high correlation between the left and right striatal SBRs and the symmetric structure of the two nigrosomes along with their surrounding regions, we propose an improved regression model with a novel loss function (Fig. 1). Furthermore, we introduce an uncertainty estimation mechanism based on the symmetry for reliable use in clinical sites.
@@ 3.3.2 Correlated Output
Furthermore, we also argue that the SBR scores of the right and left striata are also connected. Usually, dopaminergic cell loss occurs similarly on both sides. Consequently, a low SBR score on the left striatum is likely to indicate a low SBR score on the right, and vice versa. Empirically, the right and left SBRs gave a high correlation coefficient of $0.93$ based on our data. Fig. 2 presents the correlation plot of the right and left SBR values. We propose to enforce an explicit constraint regarding this relation in our symmetric regressor.
@@ 4.1 Experimental Setting
For experimental evaluation, we acquired $734$ nigral patches (right and left combined) of $367$ patients ($55.33\%$ female; $69.0\pm 9.2$ years old) from Seoul National University Bundang Hospital, South Korea. The nigral SMWI patch and SPECT SBR information of our dataset is presented in Table 2. The SMWI patches are obtained from a 3T MRI scanner. The detailed protocols of MRI scanning, SMWI image reconstruction, and DAT SPECT imaging are described in [30]. Two neuroradiologists (with 12 and 22 years of experience) annotated the nigrosome centroids in consensus for cropping the patches. The true SBR scores were automatically obtained from the SPECT images by the DATquant program by GE Healthcare.
@@
Total cases $734$ ($367$ right and $367$ left nigral patches)
@@
Data acquisition period February $2017$ December $2018$
MRI (SMWI) device Ingenia and Ingenia CX, Philips, Netherlands
Detail protocol in[30]
@@
SPECT device DATrace-123™, Samyoung Unitech, Korea
with Trionix XLT, Trionix Research Lab, USA.
Detail protocol in [30]
SBR analysis program DATquant, Xeleris 3.1, GE Healthcare, USA
SBR range Right: $[0.44,6.84]$; left: $[0.43,6.80]$
MRI/SPECT time diff. $42\pm 60$ days (median: $18$ hours)
@@ 4.2 Hyperparameter Dependency
The key hyperparameters in the proposed model are the clipping margin $\alpha$ and weight $\beta$ of the symmetric loss. We present the mean absolute error (MAE) of the SBR prediction for different values of $\alpha$ and $\beta$ in Fig. 3. Note that the optimal hyperparameters were chosen based on the validation set performance. However, the dependency plots here are presented for the test set.
Figure 3: Dependency on the clipping margin (top) and weight (bottom) of the symmetric loss. The weight axis in the bottom plot is in log scale.
The clipping margin contributes to ignoring small right-to-left differences while computing the symmetric loss. From the resultant plot in Fig. 3(top), it is apparent that a small margin gave a poor performance because it strictly penalized the objective even for the small symmetric differences. On the other hand, a large margin also gave a non-optimal performance because it ignored most of the differences including the cases where the right-to-left prediction differences were big. With the optimal clipping margin, the loss only ignored prediction differences less than $0.04$ while penalizing the objective for the others, thus improving the prediction performance.
The symmetric loss weight controls the influence of the symmetric loss over the original regression loss. We present the prediction performance for different weight values in Fig. 3(bottom). Note that the weight axis in the plot is presented in a logarithmic scale. In this experiment, we used the optimal clipping margin i.e., $\alpha=0.04$ (decided previously). A high weight puts too much focus on the symmetric loss resulting in a higher MAE. Contrarily, a low weight fails to effectively utilize the symmetric property of our input and output pairs.
@@ 4.3 SBR Prediction Performance
Following general continuous valued prediction problems, we evaluate the proposed regressor based on the MAE and root mean squared error (RMSE) between the model predictions and the true SBR values. Furthermore, we also compare Pearson’s correlation coefficient $R$, which is a widely used performance measure in the medical field. We show the improvements made by our propositions in three steps. First, we compare the proposed VGG architecture with the existing VGG-16 [23] and ResNet-18 [24] architectures. Second, we present the improvements of the paired input-output model. Finally, we compare the performance of the paired model trained with the symmetric loss. We present our comparison results in Table 3.
@@
Among the decoupled right and left SBR predictors, the proposed VGG network gave a lower error than the existing ResNet-18 and VGG-16 networks. While ResNet-18 and VGG-16 networks were originally proposed for the ImageNet dataset consisting of 2D natural images, the 3D nigral patch has a much smaller spatial dimension. The proposed network gave improved performance because it had a smaller number of layers suitable for the small input. The proposed network is identified as $VGG$-$small$ in Table 3.
@@
Compared with the individual right and left predictors, our paired model performed significantly better. In the paired model, right and left SBRs were predicted simultaneously. As a result, the model was provided with more training data on top of implicit knowledge sharing between the highly correlated right and left prediction tasks. Thereby, it gave a better generalization utilizing the symmetric property. We identify the paired model as $Symmetric$ $VGG$ in Table 3. Moreover, the symmetric loss helped train the paired model even more effectively by enforcing an additional constraint based on the output correlation in the pairs. As a result, our final model ($Symmetric$ $VGG$ + $L_{\text{sym}}$) gave a further improvement overall.

### v2

@@ Abstract
Dopamine transporter (DAT) imaging is commonly used for monitoring Parkinson’s disease (PD), where the amount of striatal DAT uptake serves as the PD severity indicator. However, DAT imaging has a high cost and the risk of radiance exposure, besides its unavailability in general clinics. Recently, MRI of the nigral region has emerged as a safer and easier alternative. This work introduces a symmetric regressor for predicting the DAT uptake amount from the nigral MRI patch. Acknowledging the lateral symmetry between the right and left nigrae, the proposed regressor incorporates a paired input-output model that concurrently predicts the DAT uptake amounts for both the right and left striata. Moreover, it employs a symmetric loss that constrains the difference between right-to-left predictions, resembling the inherent correlation in DAT uptake amounts on the two lateral sides. Additionally, this work proposes a symmetric Monte-Carlo (MC) dropout method for providing a fruitful uncertainty estimate about the DAT uptake prediction, which utilizes the above symmetry. We evaluated the proposed approach on $734$ nigral patches, demonstrating significantly improved performance of the symmetric regressor compared with the standard regressors while giving better explainability and feature representation. Furthermore, the symmetric MC dropout gave precise uncertainty ranges for model predictions with a high probability of including the true DAT uptake amount within the range.
keywords:
Brain MRI , DAT SPECT , deep learning , dopamine transporter , Monte-Carlo uncertainty , Parkinson’s disease , severity assessment
journal: *
@@ 1 Introduction
Recent studies suggested using MRI as a less invasive and readily available alternative imaging method for monitoring PD [6, 7]. In MRI-based assessment, the loss of nigral hyperintensity is generally used for PD diagnosis as it corresponds well with the low DAT uptake. Nonetheless, a quantified severity assessment is not achievable by only manual observation of such intensity. Therefore, a computerized method for predicting the DAT uptake amount (as measured in SPECT) using the nigral MRI can be substantially useful. The Deep learning-based regressor showed promising results for DAT uptake prediction [8]. However, standard deep regressors generally minimizes the mean squared error (MSE) loss for the individual samples, where the samples are considered independent and identically distributed (i.i.d). On the contrary, the DAT uptake amounts in the right and left lateral sides have high correlation. Thus, the typical regressors fail to utilize such correlation.
In this paper, we propose a symmetric deep regressor for predicting the DAT uptake amount from the MRI patch covering the nigral region. While the typical regressor follows a one-input-one-output structure with a loss function based on the mean squared error w.r.t. the known true output, our symmetric model adopts a paired input-output structure. Our regressor takes both the right and left nigral patches, simultaneously generating DAT uptake predictions for both the right and left striata. Additionally, we employ a symmetric loss that penalizes high differences between right and left predictions, addressing the symmetry between the two nigral regions and the strong correlation between DAT uptakes in the right and left striata. The proposed formulation facilitates explicit knowledge sharing between the two prediction tasks, thereby, allowing for enhanced feature learning. Figure 1 illustrates the proposed symmetric regressor model.
@@
  • 1.
    We propose a symmetric regressor with a novel loss function for the striatal DAT uptake prediction, which utilizes the lateral symmetry and correlation of the right and left nigral patch inputs and the corresponding DAT uptake outputs.
  • 2.
@@ 2.2 DAT Uptake Prediction using MRI
The use case of MRI is not limited to only PD diagnosis. Researchers have also used the nigral status in MRI to correlate with the SBR in SPECT. Bae et al. [19] compared the nigral hyperintensity observed in 3T SWI with the SPECT SBR in Parkinsonism. In their later work [20], they used a similar SWI-to-SBR comparison approach for idiopathic rapid eye movement disorder. Both of these works revealed good agreement between the SWI observation and the SBR score. Uchida et al. [21] used the mean intensity of the manually annotated striatum region in order to correlate to the SBR score. Recently, susceptibility map-weighted imaging (SMWI) [22] has been proposed for improving the image quality by providing with better contrast in neuroimaging. Nam et al. [23] utilized SMWI to enhance the visibility of nigrosome for PD diagnosis, which allowed for better identification of the hyperintensity. Bae et al. [24, 25] also compared SWI and SMWI images of the substantia nigra in Parkinsonism, and concluded that SMWI provides improved quality.
@@ 2.3 Deep Learning Based Prediction
In this work, we utilize both the right and left nigral patches to predict SBRs of both the right and left striata. Leveraging the high correlation between the left and right striatal SBRs and the symmetric structure of the two nigrosomes along with their surrounding regions, we propose an improved regression model with a novel loss function (Figure 1). Furthermore, we introduce an uncertainty estimation mechanism based on the symmetry for reliable use in clinical sites.
@@ 3.1 Data
We collected the dataset for this study from Seoul National University Bundang Hospital, South Korea. Approval of all ethical and experimental procedures and protocols was granted by the Institutional Review Board (IRB) of Seoul National University Bundang Hospital, South Korea (IRB number: B-1610-368-303). Written informed consent was obtained from all study participants.
@@ 3.3.2 Correlated Output
Furthermore, we also argue that the SBR scores of the right and left striata are also connected. Usually, dopaminergic cell loss occurs similarly on both sides. Consequently, a low SBR score on the left striatum is likely to indicate a low SBR score on the right, and vice versa. Empirically, the right and left SBRs gave a high correlation coefficient of $0.93$ based on our data. Figure 2 presents the correlation plot of the right and left SBR values. We propose to enforce an explicit constraint regarding this relation in our symmetric regressor.
@@ 4.1 Experimental Setting
For experimental evaluation, we acquired $734$ nigral patches (right and left combined) of $367$ patients ($55.33\%$ female; $69.0\pm 9.2$ years old) from Seoul National University Bundang Hospital, South Korea. Among the $367$ patiens, the number of PD patients were $287$. Note that, this dataset was previously used in Bae et al.’s work [8]. However, only left nigral patches were used in their work. The nigral SMWI patch and SPECT SBR information of our dataset is presented in Table 2. The SMWI patches are obtained from a 3T MRI scanner. The detailed protocols of MRI scanning, SMWI image reconstruction, and DAT SPECT imaging are described in [8]. Two neuroradiologists (with 12 and 22 years of experience) annotated the nigrosome centroids in consensus for cropping the patches. The true SBR scores were automatically obtained from the SPECT images by the DATquant program by GE Healthcare.
@@
Total cases $734$ ($367$ right and $367$ left
nigral patches)
@@
Data acquisition period February $2017$-December $2018$
MRI (SMWI) device Ingenia and Ingenia CX,
Philips, Netherlands. Detail
protocol in[8]
@@
SPECT device DATrace-123™, Samyoung
Unitech, Korea with Trionix
XLT, Trionix Research Lab,
USA. Detail protocol in [8]
SBR analysis program DATquant, Xeleris 3.1, GE
Healthcare, USA
SBR range Right: $[0.44,6.84]$;
left: $[0.43,6.80]$
MRI/SPECT time diff. $42\pm 60$ days
(median: $18$ hours)
@@ 4.2 Hyperparameter Dependency
The key hyperparameters in the proposed model are the clipping margin $\alpha$ and weight $\beta$ of the symmetric loss. We present the mean absolute error (MAE) of the SBR prediction for different values of $\alpha$ and $\beta$ in Figure 3. Note that the optimal hyperparameters were chosen based on the validation set performance. However, the dependency plots here are presented for the test set.
Figure 3: Dependency on the clipping margin (left) and weight (right) of the symmetric loss. The clipping margin in the left plot is normalized relative to the maximum SBR ($6.84$) of our dataset. The weight axis in the right plot is in log scale.
The clipping margin contributes to ignoring small right-to-left differences while computing the symmetric loss. From the resultant plot in Figure 3(left), it is apparent that a small margin gave a poor performance because it strictly penalized the objective even for the small symmetric differences. On the other hand, a large margin also gave a non-optimal performance because it ignored most of the differences including the cases where the right-to-left prediction differences were big. With the optimal clipping margin, the loss only ignored (normalized) prediction differences less than $0.04$ while penalizing the objective for larger differences, thus improving the prediction performance.
The symmetric loss weight controls the influence of the symmetric loss over the original regression loss. We present the prediction performance for different weight values in Figure 3(right). Note that the weight axis in the plot is presented in the logarithmic scale. In this experiment, we used the optimal clipping margin i.e., $\alpha=0.04$ (decided previously). A high weight puts too much focus on the symmetric loss resulting in a higher MAE. Contrarily, a low weight fails to effectively utilize the symmetric property of our input and output pairs.
@@ 4.3 SBR Prediction Performance
Following general continuous valued prediction problems, we evaluate the proposed regressor based on the MAE and root mean squared error (RMSE) between the model predictions and the true SBR values. Furthermore, we also compare Pearson’s correlation coefficient $R$, which is a widely used performance measure in the medical field. We present the SBR prediction performance of different regressor models including ours in Table 3. We show the improvements made by our propositions in three steps. First, we compare the proposed VGG architecture ($VGG$-$small$) with the standard VGG-16 [26] and ResNet-18 [27] architectures. Second, we present the improvements of the paired input-output model ($symmetric$ $VGG$). Finally, we compare the performance of the paired model trained with the symmetric loss ($symmetric$ $VGG$ + $L_{\text{sym}}$).
As Table 3 demonstrates, both the VGG-16 and ResNet-16 models gave a poor prediction performance, resulting an MAE above $0.915$. The $VGG$-$small$ network gave a better prediction error and correlation than the widely used VGG-16 and ResNet-18 networks, improving the error to $0.768$. The paired input-output model gave a further improvement by combining the right and left prediction tasks. The final model trained with the symmetric loss generally outperformed all the other models by dragging the prediction error down to $0.6741$.
4.4 Quality of Uncertainty
For plotting the curve, we obtain the (sharpness, CP) points by varying the interval from $0.\sigma$ to $m\sigma$, given the original interval $\sigma$ from MC dropout. Here, $m>0$ is defined such that the CP reaches $100\%$. Figure 4 presents the resultant sharpness-CP curves. This figure presents three curves given by the MC intervals of the standard VGG regressor, the MC intervals of the proposed regressor, and our symmetric MC intervals of the proposed regressor.
Both intervals of the proposed symmetric regressor had a greater AUC than the standard regressor. The interval we proposed using symmetric MC dropout gave a slightly better AUC than the usual MC dropout. For the final regressor, the curves of the symmetric MC and MC intervals seems almost similar except only in a few places. However, the proposed MC is still advantageous because high CP positions are usually required for practical application in medical sites. At high CP (e.g., above $80$% CP), the symmetric MC gave intervals with higher sharpness and precision. We discuss more on this in Section 5.4.
5 Discussion
5.1 Validity of the Proposed Model
5.1.1 CNN Model Selection
@@
As mentioned before, we did not use the standard CNN architectures for our CNN encoder in the symmetric regressor. Rather, we utilized a smaller version of the VGG-style architecture. From Table 3, we observed that the proposed VGG network ($VGG$-$small$) gave a lower error than the widely used ResNet-18[27] and VGG-16 [26] networks. While ResNet-18 and VGG-16 networks were originally proposed for the ImageNet dataset consisting of 2D natural images, the 3D nigral patch has a much smaller spatial dimension. The proposed network gave improved performance because it had a smaller number of layers suitable for the small input.
5.1.2 Advantage of the Symmetric Model
@@
Compared with the standard regressor with individual right and left predictors, our paired model performed significantly better. In the paired model, right and left SBRs were predicted simultaneously due to the lateral symmetry and correlation between the right and left nigral patches. As a result, the model was provided with more training data on top of implicit knowledge sharing between the highly correlated right and left prediction tasks. Thereby, it gave a better generalization utilizing the symmetric property. Moreover, the symmetric loss helped train the paired model even more effectively by enforcing an additional constraint based on the output correlation in the pairs. As a result, our final model ($Symmetric$ $VGG$ + $L_{\text{sym}}$) gave a further improvement overall.
The proposed symmetric regressor resulted in an average correlation of $0.7291$ ($p$ value $<0.000001$), which is a considerable increase from the standard regressor’s correlation (i.e., $0.6674$, $p$ value $<0.000001$). Note that, this resultant correlation is within the acceptable range in the medical field. The resultant MAE is also improved from $0.7675$ to $0.6741$ by the proposed method. We compared the results of $VGG$-$small$, $Symmetric$ $VGG$, and $Symmetric$ $VGG$ + $L_{\text{sym}}$ using Student’s $t$-test, which revealed that the improvements by $Symmetric$ $VGG$ (w.r.t $VGG$-$small$) and $Symmetric$ $VGG+L_{\text{sym}}$ (w.r.t $Symmetric$ $VGG$) were statistically significant ($p$ value $<0.01$ and $p$ value $<0.03$, respectively).
Table 4: Comparison with Other MRI-based DAT Uptake Analysis Works.
Prediction Correlation Binary Agreement (acc., sen., spe.)
Method; goal; data size Right Left Right Left
Mean intensity [21]; classification; $61$ $0.470$
Observation [25]; classification; $138$ ($0.652,0.587,\boldsymbol{0.897}$) ($0.667,0.596,\boldsymbol{0.932}$)
VGG regressor [8]; prediction; $367$ $0.628$ $0.707$ ($0.764,0.721,0.750$) ($0.855,0.786,0.756$)
This work; prediction; $734$ $\boldsymbol{0.716}$ $\boldsymbol{0.743}$ ($\boldsymbol{0.818},\boldsymbol{0.860},0.833$) ($\boldsymbol{0.873},\boldsymbol{0.854},0.786$)
5.2 Comparison with Previous Works
Most of the previous MRI-based SBR prediction works attempted to investigate a binary agreement between the normal/abnormal MRI and SPECT findings. For fruitful comparison, we also obtained such binary decisions from our continuous valued data. Based on the previous studies [25, 8], we used a threshold of $3.401$ (right), $3.345$ (left) to determine the normal and abnormal SBRs. Thus, we compare our work with the previous MRI-based studies in Table 4. In this table, we present the correlation for the continuous valued prediction. Moreover, we depict the accuracy, sensitivity, and specificity for the binary decision (i.e., classification of normal and abnormal SBRs).
Among the SBR prediction studies, Uchida et al. [21] with their mean striatal intensity approach could achieve a correlation of $0.470$, while the standard deep regressor of Bae et al. [8] gave a correlation of $0.667$. Following the standard regression approaches, the regressor in [8] was trained with the usual regression loss, attempting to minimize the individual sample-wise difference between the predicted and actual SBRs. Our proposed regressor utilizing the pairwise symmetric loss between the right and left SBRs could outperform the standard regressor while utilizing the same dataset as [8].
Among the SBR classification studies, the manual identification approach of Bae et al. [25] gave an average specificity of $0.915$. However, the accuracy ($0.659$) and sensitivity ($0.591$) was substantially low. Note that a good sensitivity is important in the medical field as it signifies how many abnormal findings are mistakenly categorized as being normal. The deep regressor of Bae et al. [8] gave a well-balanced specificity and sensitivity with an accuracy of $0.809$. On the other hand, our symmetric regressor gave the highest accuracy ($0.845$). In our work, the average sensitivity and specificity was $0.857$ and $0.810$, thus resulting in a significantly better SBR classification performance ($p$ value $<0.0004$, as revealed with McNemar’s test).
5.3 Explainability Analysis
We also visually investigated the extracted features by the proposed model to evaluate the discriminative nature of the learned representation of our data. We did so by mapping the extracted features into 2D space using t-SNE [35]. We plot the resultant 2D features in Figure 5. Each point in this plot is color-graded based on its true SBR value. We also plot the features extracted by the standard regressor for comparison. It is clear from the plots that the proposed model could learn comparatively more powerful feature representation that is useful for discriminating among different SBR ranges.
Figure 6 presents one example for each of the high and low SBR cases. For the low SBR case, gradients of the low SBR class were backpropagated to obtain the activation maps (Figure 6-top). On the other hand, gradients of the high SBR class were used for the high SBR case (Figure 6-bottom). Generally, the activation maps for high SBR prediction highlighted the nigrosome, parts of the substantia nigra, and parts of the cerebral peduncle. Contrarily, the activation maps for low SBR mainly highlighted the red nucleus region with a small minor focus on parts of the substantia nigra. The low SBR activation maps usually did not show any attention near the nigrosome because the nigrosome hyperintensity is lost in the low SBR cases. This coincides with previous medical research [15, 19] where nigrosome hyperintensity was used as an identifier for normal/abnormal SPECT findings.